OpenStudy (anonymous):

how do you calculate the magnitude of acceleration of an object just dropped

OpenStudy (akashdeepdeb):

It depends on what your are given! Use the 3 equations of motion! v=u+at s = ut + at^2/2 v^2 = u^2 + 2as Where v = final velocity, u = intial velocity, t = time , a is ACCELRATION, s = distance! :) Understood? :)

OpenStudy (vincent-lyon.fr):

List the forces acting on the object, then use Newton's 2nd law of motion.

OpenStudy (ookawaiioo):

Gravity is a constant acceleration. 9.8 m/s^2 F=ma Doesnt matter if you dropped it 500 feet or 10 feet. The value is the same. However speed will vary from the time released